How do you solve 2(x+1)^2=3x+7 using the quadratic equation??
Answer by stanbon(75887)

How do you solve 2(x+1)^2=3x+7 using the quadratic equation??
---
2(x^2+2x+1) = 3x+7
4x^2+x-5 = 0
---
4x^2+5x-4x-5 = 0
x(4x+5)-(4x+5) = 0
(4x+5)(x-1) = 0
x = -5/4 or x = 1
=======================
Cheers,
Stan H.
============